OpenStudy (anonymous):

5x-4 = 9x + 3 A) 1/2 B) 2 C) 1/4 D) 4

OpenStudy (anonymous):

none of your answer are correct

OpenStudy (mertsj):

none of the above

OpenStudy (anonymous):

OH sorry! it should say 5x+4 = 9x+3

OpenStudy (anonymous):

i mixed up the plus and minus

OpenStudy (anonymous):

it said that was wrong

OpenStudy (lilg132):

-4x = -1 4x = 1 4x / 4 = 1 / 4 x = ?

OpenStudy (anonymous):

so i type X?

OpenStudy (lilg132):

what do you mean

OpenStudy (anonymous):

i tried X but it said answer has to be A thru D

OpenStudy (anonymous):

x = 1/4 Which letter does 1/4 represent

OpenStudy (lilg132):

lol

OpenStudy (anonymous):

X?

OpenStudy (anonymous):

but it said that wuz wrong

OpenStudy (lilg132):

the question mark is there so you can work the last step out

OpenStudy (anonymous):

OMG Look at the list of choices you are given

OpenStudy (lilg132):

i've pretty much given the answer

OpenStudy (anonymous):

remember this is a multiple choice question

OpenStudy (anonymous):

ya but if i guess wrong i lose a point

OpenStudy (lilg132):

ok confirm it with me then

OpenStudy (lilg132):

what do you think the answer is

OpenStudy (anonymous):

A?

OpenStudy (lilg132):

why do you think its a?

OpenStudy (anonymous):

A) 1/2 B) 2 C) 1/4 D) 4 Which letter is 1/4?

OpenStudy (anonymous):

C?

OpenStudy (lilg132):

moneybird give her a chance to learn

OpenStudy (lilg132):

why do you think its c?

OpenStudy (anonymous):

c coz it says 1/4 next to it ... i get it now

OpenStudy (anonymous):

You are pretty "intellegent" as a Grade 7

OpenStudy (anonymous):

lol thanx hun :)

OpenStudy (anonymous):

it said C was ok, thanx guys :) <3 <3

OpenStudy (lilg132):

well all i can say is you've taught her well moneybird, thats the reason shes so intelligent

OpenStudy (lilg132):

your welcome although thank moneybird because i know you didnt understand my way

OpenStudy (anonymous):

i have some others?

OpenStudy (anonymous):

Remeber you can never have X as your answer since the choices are only A, B, C, and D

OpenStudy (anonymous):

ok even if X is in the problem?

OpenStudy (anonymous):

Well first you need to distinguish whether the question is a multiple choice type or not

OpenStudy (lilg132):

x represents a value x represents a number you need to work out the number

OpenStudy (lilg132):

not everyone finds that easy to do moneybird

OpenStudy (anonymous):

5t + 2t = 14 t= A) 1 B) 2 C) 3 D) 4

OpenStudy (anonymous):

i agree

OpenStudy (lilg132):

what is 5t + 2t ?

OpenStudy (anonymous):

What is 5 +2 ?

OpenStudy (anonymous):

i dont know> just need a letter i think

OpenStudy (lilg132):

hold up five fingers

OpenStudy (lilg132):

type with one hand and tell me when you've done that

OpenStudy (anonymous):